OpenStudy (anonymous):

Which of the following fractions can be used in the conversion of 1450 cm3 to the unit km3? fraction cubed: one meter over ten to the negative second power cm one mL over one cubic cm fraction cubed: one km over ten to third power cm fraction cubed: ten to the negative second power m over one cm

OpenStudy (anonymous):

\[\left( \frac{ 1m }{ 10^{-2} } \right)^{3}\] \[\frac{ 1 ml }{ 1 cm ^{3} }\] \[\left( \frac{ 1 km }{ 10^{3}cm } \right)^{3}\] \[\left( \frac{ 10m^{-2} }{ 1 cm } \right)^{3}\]